OpenStudy (anonymous):

PLEASE HELP. solve over domain -2pi

OpenStudy (anonymous):

4sinxcosx=2sin^2x 2sinxcosx=sin^2x

OpenStudy (anonymous):

sin2x=sin^2x

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

answer key has 6 answers

sam (.sam.):

Put it as quadratic equation \[4sinxcosx=2-2\cos^2x \\ \\ 2\cos^2x+4sinxcosx-2=0 \] quadratic formula \[cosx=\frac{-4\pm \sqrt{16-4(2)(-2)}}{2(2)} \\ \\ cosx=\frac{-4 \pm 4 \sqrt{2}}{4} \\ \\ cosx=-1+\sqrt{2}~|~cosx=-1-\sqrt{2}\] Find "x"

OpenStudy (anonymous):

can you solve sin2x=sin^2x or was what I was doing totally off?

sam (.sam.):

sin2x=sin^2x Using trig identity, sin2x=2sinxcosx, we get 2sinxcosx=sin^2x

OpenStudy (anonymous):

yeah i tried to get it all in sin.. how do u factor sin2x=sin^2x that more?

sam (.sam.):

\[2sinxcosx=\sin^2x \\ \\ 0=\sin^2x-2sinxcosx \\ \\ sinx(sinx-2cosx)=0 \\ \\ sinx=0~~~|~~~sinx-2cosx=0 \\ \\ tanx=2\] Then solve for sinx=0 and tanx=2

OpenStudy (anonymous):

thank youu!!!!!